OpenStudy (anonymous):

im taking an online quiz i cant mess up so what is the slope of (-5,2) and (-5,-3) what is it? i got -5 and 0 when i did y2-y1 and x2-x1 so -5/0? is that even possible help

OpenStudy (texaschic101):

when you have both x values the same, it is a vertical line. VUX -- vertical line, undefined slope, represented by x = a number......x, in this problem, will always be -5. So your slope is undefined and the equation for this is : x = -5. By the way, if you have 0 as a denominator, then it is undefined...but if you have 0 as the numerator, then it has a 0 slope

OpenStudy (texaschic101):

slope = -3 - (-2) / (-5 - (-5) slope = (-3 + 2) / (-5 + 5) slope = -1/0.......undefined slope